# Understanding differences in applying DETR to natural and medical images

### v1

@@ 4 Results
This observation aligns with the recent development of the encoder-free $\text{D}^{2}\text{ETR}$ (Lin et al., 2022), which outperforms the standard DETR model on the MS COCO dataset (Lin et al., 2014). Such findings challenge the traditional view that encoders are essential for feature transformation and multi-level feature integration within DETR models. Our results suggest that it is possible to create effective DETR architectures without encoders, avoiding significant sacrifices in performance. Additionally, a strong backbone can further reduce DETR’s reliance on the encoders, further supporting the potential for more streamlined designs.
Refer to captionRefer to caption
Refer to captionRefer to caption
Figure 4: (a) Comparison of the number of objects per image and object sizes between MS COCO and the NYU Breast Screening dataset. Compared to the COCO dataset, in the NYU Dataset object sizes vary less. The standard deviation of object sizes in COCO and NYU breast are 0.161 and 0.025 respectively. (b) An example image from the NYU Breast Dataset containing a cancerous lesion. In this data set, there are, on average, 1.10 objects per image in images that contain any objects. (c) Example images from the MS COCO dataset. Both images contain multiple objects in different sizes and some of them even overlap, adding difficulty to object detection. There are, on average, 7.33 objects per image in MS COCO dataset.
@@
Standard Deformable DETR applies the attention mechanism to four scales of feature maps in the encoder: three from the last three layers of the backbone and fourth from a convolution layer applied to final output of the backbone, as illustrated in Figure 2(b). Previous findings highlight that using multi-scale feature fusion improves detection performance by 1.3 on the MS COCO dataset and also on other datasets (He et al., 2017; Zhou et al., 2021; Zeng et al., 2022), our results in Table 3 indicate that comparable performance can be achieved using only the output of the last layer of backbone. This suggests that multi-scale feature fusion may not be crucial for detecting abnormalities in medical images.
@@
The characteristics of our dataset likely explain this finding. Firstly, our dataset predominantly consists of images containing a single object. Secondly, the sizes of these objects are relatively uniform. These characteristics contrast with natural image datasets such as MS COCO, where object sizes vary widely due to perspective, distance from the camera, and the inherent size differences between object classes, as shown in Figure 4(c). Multi-scale feature fusion allows models to be sensitive to this size variability by utilizing feature maps at different scales, each of which captures details at different resolutions. However, our dataset is more homogeneous, with images containing single objects of similar sizes (Figure 4(b). For such data, the benefits of multi-scale feature fusion are less pronounced. Figure 4(a) contrasts the object size distribution between the two datasets. In our dataset, the majority of images contain only a single object and most objects take only approximately $1\%$ of the entire image, whereas MS COCO dataset exhibits a broader variation in both the sizes of objects and the number of objects per image. Consequently, in a homogeneous dataset, the additional complexity of multi-scale feature fusion may not translate into better performance.
@@
Table 3: The performance of Deformable DETRs using different combinations of feature levels. The standard Deformable DETR uses the 4 levels of feature maps from the backbone (three from the backbone and the fourth layer from a convolution layer applied to final output of the backbone). Using only the 3rd feature level, the last layer of the backbone, achieves on-par performance with using multi-level feature fusion. All experiments are conducted on model with 3 encoder layers (cf. Table 2).
Feature Levels $\text{FAUC}^{1}_{10}$ $\text{AP}_{10}$ $\text{AP}_{10,50}$ params GFLOPs
$1,2,3,4$ (standard) $0.688\pm 0.012$ $0.677\pm 0.011$ $0.470\pm 0.019$ $38.3$ $836$
$1$ $0.637\pm 0.012$ $0.627\pm 0.009$ $0.434\pm 0.009$ $35.5$ $734$
$2$ $0.680\pm 0.005$ $0.666\pm 0.009$ $0.467\pm 0.016$ $35.6$ $570$
$3$ $0.688\pm 0.004$ $0.675\pm 0.007$ $0.475\pm 0.006$ $35.7$ $528$
@@
(a) $\text{FAUC}^{1}_{10}$
(b) $\text{AP}_{10}$
(c) $\text{AP}_{10,50}$
(d) $L$ and $L_{top10}$
Figure 5: (a)-(c) The detection performance of Deformable DETR with varying number of object queries. Detection performance improves as the number of object queries increases from 5 to 100. However, performance declines when the number exceeds 100. (d) The localization performance $L$ continues to improve as the number of object query increases. The classification performance $L_{\mathrm{top10}}$ worsens as number of object queries surpass 100, indicating an increase in false positives.
Figure 5(a)-(c) depicts the impact of increasing the number of object queries from 5 to 500 on three detection metrics. Increasing the number of object queries from $5$ to $100$ consistently improves detection performance. However, exceeding this number does not bring additional benefits and may even slightly hurt performance. The benefit of more object queries for object localization is evident in Figure 5(d), which shows improved localization performance $L$ (as defined in Equation 1) as the number of queries increases. However, better localization does not necessarily equate to improved overall detection performance, as indicated by $L_{\mathrm{top10}}$ (cf. Equation 2), the proportion of correctly localized objects that are also among top $10$ classification scores. Although the model correctly identifies more ground truth objects, it assigns them low classification scores, resulting in fewer objects ranked among the top 10 bounding boxes.
We hypothesize that having more object queries increases the chances of localizing false positives. More object queries expand model’s search space, making it more sensitive to variations in subtle patterns or textures that may resemble the characteristics of positive objects. These false positives cause true positives to be ranked lower. In summary, while more object queries improves the chances of detecting an object, they also introduce the risk of misidentifying irrelevant entities as objects of interest, thus adversely impacting the model’s overall performance. This phenomenon can be more severe in medical imaging datasets, where there are significantly fewer objects (usually a single object) within an image, increasing the likelihood of false positives overwhelming the true positives.
@@
We implemented two techniques commonly adopted in DETR family using the simplified model with $3$ encoder layers and $1$ feature map: query initialization method and iterative bounding box refinement (IBBF). Surprisingly, empirical results indicate that neither technique significantly improved detection performance across all three metrics, as detailed in Table 4. Following previous analysis, we examine the model’s localization and classification performance separately by evaluating on $L$ and $L_{\mathrm{top10}}$. We found that while these techniques enhanced localization performance, they adversely affected classification performance, as depicted in Figure 6.
We visualize training and validation losses for localization and classification separately in Figure 7. Localization losses include the IoU loss and bounding boxes regression loss, while classification loss is the BCE loss on all prediction boxes. The figure reveals that the model equipped with either decoding techniques tends to overfit the training data more than the model without them, particularly in classification loss, as illustrated in Figure 7. This overfitting is likely the reason behind the observed decline in classification performance. We hypothesize that this is due to the significantly smaller number of positive objects in our dataset.
Table 4: Deformable DETR with different query initialization methods and IBBF. Neither of the techniques improve detection performance across all three metrics. All experiments are conducted on model with $3$ encoder layers and without multi-scale feature fusion technique.
Refinement Query Initial. $\text{FAUC}^{1}_{10}$ $\text{AP}_{10}$ $\text{AP}_{10,50}$
Static $0.688\pm 0.004$ $0.677\pm 0.011$ $0.470\pm 0.019$
Pure $0.678\pm 0.007$ $0.668\pm 0.011$ $0.474\pm 0.018$
Mixed $0.670\pm 0.012$ $0.633\pm 0.010$ $0.443\pm 0.014$
$\checkmark$ Static $0.680\pm 0.010$ $0.678\pm 0.012$ $0.470\pm 0.009$
$\checkmark$ Pure $0.670\pm 0.010$ $0.670\pm 0.019$ $0.468\pm 0.013$
Figure 6: The x-axis of the figure indicates the decoding technique used in the model listed in Table 4. All the decoding techniques improve the localization performance $L$. However, the classification performance $L_{\mathrm{top10}}$ does not improve accordingly, indicating an increase in false positives.
Cases Visualization
@@
Finally, to better understand which test examples are easy and hard for our models, we visualized a few exams along with their classification scores. Figures 8 and 9 showcase images where the model assigned cancerous objects high malignant scores (scores $\geq 0.8$) and low scores (scores $\leq 0.1$), respectively. We observed that the model correctly localizes abnormal objects in all images. However, it tends to assign high scores to high-density masses featuring non-circumscribed, irregular, or indistinct borders, which are typically indicative of malignancy to the human eye. Conversely, the model usually assigns low scores to low-density masses with circumscribed borders, which can easily be confused with benign cases (Lee et al., 2018).
Figure 7: Training and validation loss of DETR models. The left panel depicts the classification loss and the right panel depicts the localization loss. Each line represents the average of five models with different random seeds. It clear that the models without any decoding techniques (the blue line) overfit significantly less to the training data than other models with decoding techniques (other color lines).

### v2

@@ 4 Results
These results suggest that the encoder can be shallower in DETR, regardless of whether the backbone is pretrained. When a powerful, domain-adapted backbone is available, the encoder can be removed with minimal impact on performance. This observation aligns with the recent development of the encoder-free $\mathrm{D}^{2}\mathrm{ETR}$ (Lin et al., 2022), which outperforms the standard DETR model on the MS COCO dataset (Lin et al., 2014). Together, these insights challenge the conventional view that encoders are essential for feature transformation and multi-level feature integration within DETR models. Our results suggest that effective DETR-based detection can be achieved without encoders, particularly when paired with powerful backbones, offering a promising path toward more efficient and streamlined model designs.
Table 2: Varying the number of encoder layers in Deformable DETR. The standard Deformable DETR has $6$ encoder layers. For both datasets, we do not observe any significant performance drop when using fewer encoder layers.
Dataset backbone #encoder $\mathrm{FAUC}^{1}_{10}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pm\mathrm{SD}}$ $\mathrm{AP}_{10}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pm\mathrm{SD}}$ $\mathrm{AP}_{10,50}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pm\mathrm{SD}}$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L\pm\mathrm{SD}}$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L_{top10}\pm\mathrm{SD}}$params GFLOPs layers NYU ResNet50 00 $0.643\pm 0.020$ $0.619\pm 0.010$ $0.421\pm 0.025$ $0.910\pm 0.013$ $0.819\pm 0.010$ $35.4$ $536$ ResNet50 $1$ $0.656\pm 0.011$ $0.624\pm 0.016$ $0.439\pm 0.005$ $0.909\pm 0.005$ $0.814\pm 0.011$ $36.2$ $624$ ResNet50 $3$ $0.655\pm 0.016$ $0.627\pm 0.013$ $0.439\pm 0.015$ $0.910\pm 0.009$ $0.818\pm 0.014$ $37.7$ $801$ ResNet50 $6$ $0.657\pm 0.009$ $0.626\pm 0.009$ $0.436\pm 0.011$ $0.910\pm 0.015$ $0.820\pm 0.02$ $40.0$ $1067$ NYU Swin-T 00 $0.681\pm 0.013$ $0.662\pm 0.013$ $0.458\pm 0.023$ $0.923\pm 0.006$ $0.843\pm 0.008$ $35.9$ $570$ Swin-T $1$ $0.684\pm 0.009$ $0.672\pm 0.011$ $0.463\pm 0.014$ $0.918\pm 0.007$ $0.841\pm 0.006$ $36.7$ $659$ Swin-T $3$ $0.688\pm 0.012$ $0.677\pm 0.011$ $0.470\pm 0.019$ $0.918\pm 0.005$ $0.855\pm 0.011$ $38.3$ $836$ Swin-T $6$ $0.689\pm 0.008$ $0.669\pm 0.012$ $0.464\pm 0.019$ $0.920\pm 0.005$ $0.856\pm 0.009$ $40.5$ $1102$ LUNA Swin-T 00 $0.011\pm 0.011$ $0.001\pm 0.000$ $0.0\pm 0$ 0.0 $0.433\pm 0.085$ $0.089\pm 0.022$ $35.9$ $1078$ Swin-T $1$ $0.538\pm 0.013$ $0.390\pm 0.005$ $0.296\pm 0.014$ $0.966\pm 0.002$ $0.680\pm 0.016$ $36.7$ $1225$ Swin-T $3$ $0.536\pm 0.010$ $0.399\pm 0.003$ $0.292\pm 0.008$ $0.966\pm 0.008$ $0.672\pm 0.008$ $38.3$ $1521$ Swin-T $6$ $0.537\pm 0.007$ $0.392\pm 0.013$ $0.295\pm 0.005$ $0.959\pm 0.002$ $0.663\pm 0.011$ 40.5 $1966$
@@
Standard Deformable DETR uses four feature maps of different scales in the encoder: three from the last three layers of the backbone and a fourth from a convolution applied to the backbone’s final output, (Figure 3(b)). Previous work show that multi-scale feature fusion improves detection performance on the MS COCO dataset as well as on other datasets (He et al., 2017; Zhou et al., 2021; Zeng et al., 2022). However, our results in Table 3 indicate that comparable performance can be achieved using only a single feature map of the backbone. This suggests that multi-scale feature fusion may not be necessary for detecting abnormalities in medical images.
@@
The characteristics of medical datasets likely explain this finding. The objects in natural image datasets, such as MS COCO, show a high variability in scale and quantity due to perspective, camera distance, and the inherent size differences between object classes (Figure 2). Multi-scale fusion benefits such settings by enabling the model to attend to features at different resolutions, capturing objects of varying sizes more effectively. However, in medical datasets like NYU Breast and LUNA16, most images contain a single object and the sizes of these objects are relatively uniform (Figure 6). This contrasts to the MS COCO dataset, showing a broader variation in both the sizes of objects and the number of objects per image. For such medical datasets, the benefits of multi-scale feature fusion are less pronounced. Consequently, in a homogeneous dataset, the additional complexity of multi-scale feature fusion may not translate into better performance.
Notably, in LUNA16, we observed that using the last feature level resulted in a performance drop. This is likely due to the extremely small object size in LUNA16, where all nodules occupy on average 0.6% of the image area (Figure 6 (b)). The last-layer feature map has too low a spatial resolution (i.e. downsized to $16\times 16$) to preserve the fine-grained detail necessary for detecting such small objects. This highlights that while multi-scale fusion may not be generally required in medical imaging, selecting an appropriate single feature level, especially one with sufficient spatial resolution, is still critical for detecting very small targets.
Figure 6: Comparison of object count per image and object size variability across MS COCO, NYU Breast, and LUNA16 datasets. In contrast to the MS COCO dataset, the medical imaging datasets, NYU Breast and LUNA16, contain fewer objects per image and show significantly less variation in object size. The standard deviation of normalized object sizes is 0.161 for COCO, 0.025 for NYU Breast, and 0.001 for LUNA16.
@@
Table 3: The performance of Deformable DETRs using different combinations of feature levels. The standard Deformable DETR uses all 4 levels of feature maps from the backbone. Using only the 3rd level feature map for breast NYU and 2nd level feature map for LUNA16, achieves on-par with or even better performance than using multi-level feature fusion.
Dataset Feature Levels $\mathrm{FAUC}^{1}_{10}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pm\mathrm{SD}}$ $\mathrm{AP}_{10}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pm\mathrm{SD}}$ $\mathrm{AP}_{10,50}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pm\mathrm{SD}}$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L\pm\mathrm{SD}}$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L_{top10}\pm\mathrm{SD}}$ params GFLOPs NYU $1,2,3,4$ (standard) $0.688\pm 0.012$ $0.677\pm 0.011$ $0.470\pm 0.019$ $0.918\pm 0.005$ $0.855\pm 0.011$ $38.3$ $836$ $1$ $0.637\pm 0.012$ $0.627\pm 0.009$ $0.434\pm 0.009$ $0.924\pm 0.007$ $0.849\pm 0.016$ $35.5$ $734$ $2$ $0.680\pm 0.005$ $0.666\pm 0.009$ $0.467\pm 0.016$ $0.917\pm 0.005$ $0.851\pm 0.012$ $35.6$ $570$ $3$ $0.688\pm 0.004$ $0.675\pm 0.007$ $0.475\pm 0.006$ $0.915\pm 0.010$ $0.858\pm 0.009$ $35.7$ $528$ LUNA $1,2,3,4$ (standard) $0.538\pm 0.013$ $0.390\pm 0.005$ $0.296\pm 0.014$ $0.966\pm 0.002$ $0.680\pm 0.016$ $36.7$ $1225$ $1$ $0.540\pm 0.007$ $0.378\pm 0.022$ $0.267\pm 0.020$ $0.961\pm 0.004$ $0.674\pm 0.022$ $34.2$ $1141$ $2$ $0.548\pm 0.017$ $0.403\pm 0.006$ $0.309\pm 0.008$ $0.965\pm 0.003$ $0.668\pm 0.007$ $34.3$ $1018$ $3$ $0.491\pm 0.015$ $0.290\pm 0.025$ $0.142\pm 0.019$ $0.959\pm 0.005$ $0.637\pm 0.010$ $34.4$ $987$
(a) $\mathrm{FAUC}^{1}_{10}$
(b) $\mathrm{AP}_{10}$
(c) $\mathrm{AP}_{10,50}$
(d) $L$ and $L_{top10}$
Figure 7: (a–c) Detection performance of Deformable DETR across varying numbers of object queries on the NYU Breast and LUNA16 datasets. Performance improves as the number of queries increases from 5 to 100 but hit plateau when the number exceeds 100. (d) Localization performance ($L$) continues to improve with more queries, while classification performance ($L_{\mathrm{top10}}$) drops beyond 100 queries, suggesting an increase in false positives that displace true positives in the top-ranked predictions.
@@
Figure 7(a)-(c) illustrates the impact of increasing the number of object queries from 5 to 800 on detection performance across both the NYU Breast and LUNA16 datasets. Increasing the number of object queries from $5$ to $100$ consistently improves the detection performance. However, further increasing the number of queries beyond 100 results in diminishing returns and even a slight decline in performance. This pattern is consistent on both datasets, although more obvious on LUNA16.
To better understand this behavior, we examined the performance on localization $L$ (cf. Equation 1) and classification $L_{\mathrm{top10}}$ (cf. Equation 2) separately in Figure 7(d). Localization performance ($L$) continues to improve with more object queries, indicating an enhanced ability to correctly localize objects. However, the classification performance ($L_{\mathrm{top10}}$), which measures how many correctly localized boxes rank among the top 10 predictions by classification score, declines beyond 100 queries. This suggests that while more queries increase the likelihood of finding true objects, they also introduce additional false positives that dilute the ranking of true positives.
We hypothesize that having more object queries increases the chances of localizing false positives. More object queries expand the model’s search space, making it more sensitive to subtle features or noise that resemble the characteristics of true objects. This can lead to more false positives being assigned high classification scores, pushing true positives lower in the ranked predictions. This issue is especially relevant in medical imaging, where images typically contain only one or very few objects of interest. In such sparse-object settings, the increased false positive rate from excessive queries can outweigh the benefits of improved localization.
@@
We evaluated two widely used decoding techniques in the DETR family, query initialization and iterative bounding box refinement (IBBR), using a simplified model with design choices achieved through previous results. As shown in Table 4, neither technique significantly improved detection performance across $\mathrm{FAUC}^{1}_{10}$, $\mathrm{AP}_{10}$, or $\mathrm{AP}_{10,50}$ for both datasets.
To better understand this outcome, we separately analyzed localization and classification performance using $L$ and $L_{\mathrm{top10}}$. We found that while these techniques improved localization performance, they adversely affected classification performance. Figure 8 shows training and validation losses for localization (IoU and box regression) and classification (binary cross-entropy). Models equipped with query initialization or IBBR display stronger overfitting, especially in classification loss, compared to the baseline model without these techniques. We hypothesize that this overfitting is due to the limited number of positive objects in our datasets. As the model becomes more effective at localizing regions of interest, it may focus too narrowly on those few positive examples, leading to memorization rather than learning generalizable features. This reduces the model’s ability to accurately distinguish between subtle classes, ultimately weakening the classification performance.
Table 4: Impact of query initialization methods and iterative bounding box refinement (IBBR) on Deformable DETR performance. Detection performance is evaluated on models with the best configuration based on previous experiments for each dataset. Neither query initialization strategies (static, pure, or mixed) nor IBBR consistently improve performance across the three main detection metrics ($\mathrm{FAUC}^{1}_{10}$, $\mathrm{AP}_{10}$, $\mathrm{AP}_{10,50}$). While some configurations improve localization ($L$), classification performance ($L_{\mathrm{top10}}$) generally declines.
Dataset Refinement Query Initial. $\mathrm{FAUC}^{1}_{10}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pm\mathrm{SD}}$ $\mathrm{AP}_{10}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pm\mathrm{SD}}$ $\mathrm{AP}_{10,50}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pm\mathrm{SD}}$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L\pm\mathrm{SD}}$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L_{top10}\pm\mathrm{SD}}$ NYU Static $0.688\pm 0.004$ $0.675\pm 0.011$ $0.475\pm 0.019$ $0.915\pm 0.010$ $0.858\pm 0.009$ Pure $0.678\pm 0.007$ $0.668\pm 0.011$ $0.474\pm 0.018$ $0.928\pm 0.011$ $0.833\pm 0.015$ Mixed $0.670\pm 0.012$ $0.633\pm 0.010$ $0.443\pm 0.014$ $0.936\pm 0.005$ $0.834\pm 0.012$checkmark Static $0.680\pm 0.010$ $0.678\pm 0.012$ $0.470\pm 0.009$ $0.931\pm 0.003$ $0.848\pm 0.009$checkmark Pure $0.670\pm 0.010$ $0.670\pm 0.019$ $0.468\pm 0.013$ $0.957\pm 0.008$ $0.838\pm 0.007$ LUNA Static $0.548\pm 0.017$ $0.403\pm 0.006$ $0.309\pm 0.008$ $0.965\pm 0.003$ $0.668\pm 0.007$ Pure $0.520\pm 0.007$ $0.396\pm 0.006$ $0.298\pm 0.008$ $0.965\pm 0.006$ $0.651\pm 0.008$ Mixed $0.500\pm 0.018$ $0.383\pm 0.007$ $0.291\pm 0.013$ $0.967\pm 0.006$ $0.636\pm 0.022$checkmark Static $0.516\pm 0.013$ $0.393\pm 0.017$ $0.303\pm 0.022$ $0.969\pm 0.014$ $0.639\pm 0.004$checkmark Pure $0.521\pm 0.015$ $0.399\pm 0.003$ $0.300\pm 0.003$ $0.965\pm 0.007$ $0.635\pm 0.005$
(a) NYU Breast
(b) LUNA16
Figure 8: Training and validation losses of Deformable DETR models with and without decoding techniques for (a) NYU Breast dataset and (b) LUNA16 dataset. Each plot shows the average loss over five runs with different random seeds. The left panels display classification loss (binary cross-entropy), and the right panels show localization loss (GIoU + bounding box regression). Models without decoding techniques (blue lines) consistently show less overfitting, especially in classification loss, compared to models using query initialization or iterative bounding box refinement.
Cases Visualization on NYU Breast
@@
Finally, to better understand how DETR models make predictions, we visualized a few exams along with their classification scores on NYU Breast dataset. Figures 9 and 10 show images in which the model assigned cancerous objects high malignant scores (scores $\geq 0.8$) and low scores (scores $\leq 0.1$), respectively. We observed that the model correctly localizes abnormal objects in all images. However, it tends to assign high scores to high-density masses featuring non-circumscribed, irregular, or indistinct borders, which are typically indicative of malignancy to the human eye. In contrast, the model usually assigns low scores to low-density masses with circumscribed borders, which can easily be confused with benign cases (Lee et al., 2018).